Full explanation would be appreciated!The demand for books is: Qd 120-P The supply of books is: OS = 5P a) Find equilibrium price and quantity b) If P = $15, which of the following is true? A) There is a surplus equal to 30 B) There is a shortage equal to 30 C) There is a surplus, but it is impossible to determine how large D) There is a shortage, but it is impossible to determine how large 32
Answer #1

a) At equilibrium prices, demand and supply are equal. I.e. Qd = QS


Qd = 120-p Qs = 5p


= 120 = 5p + p


= 120 = 6P


P = 120/6. The equilibrium price of the commodity will be 20.


At the equilibrium price, a good demand is 120 -p = 120 -20 or 100.


b) When P is $15, the quantity supplied is 5(15) = 75, and the quantity demanded is Qd = 120 -15 = 105.


Demand is greater than supply and there will be a shortage of 30. The answer is "B".


answered by: Gavin